OpenStudy (anonymous):

solve for X... e^x-1-2.3 (this is a logarithm section of the book)

OpenStudy (anonymous):

\[e^{x-1}-2.3\]?

OpenStudy (anonymous):

oops, that should read =2.3

OpenStudy (anonymous):

\[e ^{x-1}=2.3\]

OpenStudy (anonymous):

Take ln of both side ln e^(x-1)=ln (2.3) x-1=ln(2.3)

OpenStudy (anonymous):

** ln and e cancel, leaving us with just (x-1)

OpenStudy (anonymous):

x=ln(2.3)+1

OpenStudy (anonymous):

wait a sec... ln and e cancel???

OpenStudy (anonymous):

yes, ln e = 1

OpenStudy (anonymous):

I see... ok, thanks!